OpenStudy (anonymous):

Solve for x: 2.00*10^5= x(2x)^2

OpenStudy (anonymous):

x*4*x^2=2*10^5 4*x^3=2*10^5 x^3=1/2 *10^5 x^3=10/2 *10^4 x^3=50*10^3 x=10*(50)^1/3